A car traveling at 60 km/h overtakes another car traveling at 42 km/h. Assuming each car to be 5m long find the time taken during the overtake and the total ROAD distance used for the overtake.

Grade:11

1 Answers

Arun
6 years ago

For the car with 60kmph to overtake the first car it should travel a relative distance of 10 m

 

(5m for catching up and 5m for getting clear ahead)

 

therefore Srel = 10 m

 

      Vrel = 60-42 = 18km/hr = 5 m/s

 

      Relative acceleration = 0 .since both are moving with constant velocities.

 

Now using relative motion,           S = Vrel × t

 

therefore time ,t = S / Vrel = 10/5 = 2s

 

Distance is   60 × 5/18 × 2 = 100/3 = 33.33 m

hence total road distance used for overtake = 5 + 33.33 = 38.33 m